suppose that a box has a base with a width of x, a length of x+1 and a hieght of 2 inches. It is cut from a rectangular sheet of materrial with an area of 182 inches2 find the dimension of the box.
Answer by ankor@dixie-net.com(22740)

a box has a base with a width of x, a length of x+1 and a height of 2 inches.
It is cut from a rectangular sheet of material with an area of 182 inches2
find the dimension of the box.
:
From the height, (2 inches) the sheet dimensions will be:
(Add 4 inches to each box dimension):
(x+4) * (x+5) = 182
FOIL
x^2 + 5x + 4x + 20 = 182
:
x^2 + 9x + 20 - 182 = 0
:
x^2 + 9x - 162 = 0
:
Factors to:
(x + 18)(x - 9) = 0
;
Positive solution
x = 9 inches is the width of the box
and
9 + 1 = 10 inches is the length
;
;
Check solution by finding the area of the sheet
(9+4) * (10+4) = 182